# VerifyBench: A Systematic Benchmark for Evaluating Reasoning Verifiers Across Domains

### v2

@@ Abstract
Large language models (LLMs) increasingly rely on reinforcement learning (RL) to enhance their reasoning capabilities through feedback. A critical challenge is verifying the consistency of model-generated responses and reference answers, since these responses are often lengthy, diverse, and nuanced. Rule-based verifiers struggle with complexity, prompting the use of model-based verifiers. However, specialized verifiers lack flexibility, while general LLM judges can be inconsistent. Existing research primarily focuses on building better verifiers, yet a systematic evaluation of different types of verifiers’ performance across domains remains lacking, severely constraining the reliable development of Reinforcement Learning with Verifiable Reward (RLVR). To address this, we propose VerifyBench–a cross-domain comprehensive benchmark for systematically evaluating verifiers. We construct 4,000 expert-level questions covering mathematics, physics, chemistry, and biology. Each question is equipped with reference answers and diverse responses. The reliability of the evaluation is ensured through a rigorous annotation process conducted by a multidisciplinary expert team. We design a four-dimensional experimental framework to comprehensively compare the performance boundaries of specialized verifiers and general LLMs under combined conditions of extracted answers vs. complete responses, and short vs. long outputs. Our evaluation uncovers fundamental trade-offs in verifiers: while specialized verifiers achieve leading accuracy (the best model reaching 96.48% in chemistry), they exhibit deficiencies in recall; general models show stronger inclusivity but unstable precision. More importantly, we discover verifiers’ high sensitivity to input structure and inherent limitations in cross-domain generalization, providing critical insights into the bottlenecks of current verifier technology.
@@ Introduction
To fill this critical void, we introduce VerifyBench: a cross-domain comprehensive benchmark for systematic evaluation of verifiers. VerifyBench is constructed from 4,000 expert-level questions spanning mathematics, physics, chemistry, and biology. Each question includes reliable reference answers and diverse Chain-of-Thought (CoT) (Ma et al. 2025b; Wei et al. 2022) responses. Gold-standard judgment labels are established through a rigorous, fine-grained human annotation process by a multidisciplinary expert team, ensuring unparalleled reliability. This enables systematic analysis of verifier behavior across a four-dimensional experimental framework, varying input granularity (boxed final answers vs. full reasoning traces) and output constraints.
@@
We introduce VerifyBench, a novel, multidisciplinary benchmark of 4,000 expert-level questions across mathematics, physics, chemistry, and biology, featuring fine-grained human annotations and diverse CoT responses from state-of-the-art LLM, specifically designed for systematic verifier evaluation.
@@
We curate a total of 4,000 high-quality, expert-level questions across four disciplines: mathematics, physics, chemistry, and biology (1,000 per subject). These questions are designed to be challenging, incorporating rich domain-specific background, requiring advanced reasoning capabilities from LLMs, and allowing for open-ended, expressive answer spaces—thereby increasing benchmark complexity. Our definition of “expert-level” questions refers to problems typically encountered in university-level coursework or national/international academic competitions, characterized by their requirement for multi-step reasoning and synthesis of multiple concepts.
@@
Drawing from both domestic and international sources, team members manually select problems that align with our “expert-level” definition. These questions are chosen to ensure broad topical coverage, high cognitive demand, and diversity in conceptual, theoretical, and applied reasoning, always paired with correct and clear reference answers. A detailed description of our question selection criteria and source identification process is provided in Appendix.
Each question undergoes 2 rounds of manual review. Cross-validation is conducted by graduate-level (Master’s and PhD) reviewers to identify redundant knowledge points, overly simplistic phrasing, ambiguous wording, or irrelevant content. This careful review process ensures that the questions and reference answers are professional, coherent, and free from noise. Further specifics on our review protocols and reviewer qualifications can be found in Appendix.
@@
Tolerances are defined for spelling, punctuation, and formatting. In long-chain reasoning, correctness of logic and final answer are prioritized over surface-level language differences.
@@
This section offers a detailed summary of VerifyBench, highlighting its scale, composition, and core characteristics across disciplines. The benchmark contains 4,000 questions, each paired with a reference answer and a model-generated response, intended to support comprehensive and rigorous verifier evaluation. To ensure balanced representation, we include 1,000 questions from each of four major scientific domains: mathematics, physics, chemistry, and biology. Key statistics are summarized in Table 1.
@@
Total Questions | 4,000 |
Questions per Discipline | 1,000 |
@@
Total Annotated Instances | 4,000 |
@@
Verifier Math Chemistry Biology Physics Overall Trained Verifier xVerify-0.5B-I 79.60%\64.92% 94.77%\96.50% 86.75%\92.86% 92.48%\94.86% 88.77%\88.66% xVerify-3B-Ia 81.28%\56.91% 94.77%\96.21% 89.56%\90.18% 91.57%\92.24% 89.24%\85.35% xVerify-8B-I 81.58%\55.80% 96.39%\96.50% 87.95%\91.07% 92.68%\92.69% 90.03%\85.47% xVerify-9B-C 83.03%\54.14% 96.28%\94.46% 89.96%\91.07% 93.78%\92.81% 90.96%\84.76% general-verify - - - - - R1-Distill-Verifier-1.5B - - - - - General LLM as Judge Qwen2.5-7B-Instruct 78.18%\91.99% 89.45%\82.22% 68.67%\83.04% 86.66%\98.17% 83.52%\93.74% Qwen2.5-14B-Instruct 78.40%\90.61% 90.05%\82.80% 54.62%\83.04% 91.28%\96.46% 84.10%\91.55% Qwen2.5-32B-Instruct 80.98%\76.24% 85.63%\61.81% 61.45%\66.96% 95.39%\97.26% 85.34%\83.58% Qwen3-8B/14B/32B - - - - - DS-R1-Distill-Qwen-7B/14B/32B - - - - -
VerifyBench emphasizes long-form, multi-step reasoning, representing a substantial challenge for existing verifiers. The high average token count for both responses (Table 1) reflects the benchmark’s focus on detailed, explanatory solutions and verbose model outputs. Our two-stage human annotation pipeline ensures high-quality binary labels, with inter-annotator agreement (IAA) (Artstein 2017) scores ranging from 0.88 to 0.92, confirming annotation reliability.
@@
To simulate realistic response patterns of reasoning-capable language models, we employ QwQ-32B (Team 2025) to generate detailed CoT responses for all 4,000 multi-subject questions in VerifyBench. We use the official recommended hyperparameters: temperature 0.6, TopP 0.95, MinP 0, TopK 40, and no repetition penalty. The maximum response length for QwQ-32B is set to 32,768 tokens.
@@
We evaluate two broad categories of LLMs as verifiers. The first category comprises general-purpose open-source LLMs, including the Qwen2.5-Instruct series (7B, 14B, 32B) (Qwen et al. 2025), the Qwen3 series (8B, 14B, 32B) (Yang et al. 2025), and the DeepSeek-R1-Distill-Qwen series (7B, 14B, 32B) (Guo et al. 2025). The second category consists of specialized verifiers that have been trained on large-scale verification datasets, such as the xVerify series (0.5B, 3B, 8B, 9B) (Chen et al. 2025), general_verifier(Ma et al. 2025a), and R1-Distill-Verifier-1.5B (Huang et al. 2025).
Among these, models from the xVerify and Qwen2.5-Instruct series directly output binary verification decisions (e.g., “correct” or “incorrect”). In contrast, generative models—such as the Qwen3 and DeepSeek-R1-Distill-Qwen series, as well as general_verifier and R1-Distill-Verifier-1.5B—produce reasoning traces alongside their final judgments.
@@
We consider the following four experimental settings:
@@
Verifier Math Chemistry Biology Physics Overall Trained Verifier xVerify-0.5B-I 79.38%\54.14% 94.67%\94.46% 87.15%\91.07% 92.38%\92.81% 88.64%\84.76% xVerify-3B-Ia 81.18%\56.91% 95.03%\96.21% 88.76%\90.18% 91.88%\92.24% 89.38%\85.35% xVerify-8B-I 81.58%\55.80% 96.28%\96.50% 89.16%\91.07% 92.58%\92.69% 90.06%\85.47% xVerify-9B-C 82.78%\64.92% 96.18%\96.50% 89.96%\92.86% 93.98%\94.86% 90.86%\88.66% general-verify 68.77%\88.12% 75.13%\88.63% 73.09%\85.71% 94.32%\97.72% 79.01%\93.03% R1-Distill-Verifier-1.5B 76.18%\81.22% 80.71%\86.30% 77.91%\78.57% 88.77%\89.50% 81.91%\86.36% General LLM as Judge Qwen2.5-7B-Instruct 77.88%\47.51% 89.45%\93.59% 54.62%\29.46% 86.66%\86.64% 82.35%\75.90% Qwen2.5-14B-Instruct 78.08%\59.12% 90.15%\86.01% 62.25%\34.82% 91.67%\92.47% 84.75%\80.21% Qwen2.5-32B-Instruct 81.08%\71.27% 85.28%\96.50% 62.25%\48.21% 95.39%\97.37% 81.20%\88.36% Qwen3-8B 70.77%\74.31% 88.04%\93.88% 81.53%\83.04% 95.09%\97.37% 84.38%\90.79% Qwen3-14B 85.39%\80.11% 92.61%\95.92% 84.34%\92.86% 96.99%\98.52% 91.11%\93.68% Qwen3-32B 74.67%\70.72% 83.82%\93.00% 83.40%\91.96% 95.89%\97.15% 84.69%\90.31% DS-R1-Distill-Qwen-7B 64.66%\74.59% 75.38%\85.42% 74.80%\77.68% 91.78%\97.03% 77.07%\88.72% DS-R1-Distill-Qwen-14B 76.18%\78.73% 81.06%\85.13% 68.67%\64.29% 95.69%\97.26% 83.02%\88.66% DS-R1-Distill-Qwen-32B 72.37%\78.73% 79.10%\96.50% 72.29%\85.71% 96.59%\99.43% 81.85%\93.50%
@@
Here, the verifier receives the full reasoning process as input but remains limited in output length. This setting challenges the model to efficiently extract conclusions from complex context. Specialized verifiers again show dominant performance, particularly in accuracy. For instance, xVerify-9B-C yields the best overall accuracy (90.96%) and leads in most subjects. While general LLMs such as Qwen2.5-7B-Instruct demonstrate high recall (e.g., 91.99% in math), they often trade precision for inclusiveness. This highlights a key contrast: general LLMs can tolerate diverse answer forms but struggle with precision. The details are shown in Table 2.
@@
From a system perspective, verifiers should ideally produce direct, structured outputs (e.g., “Correct”/“Incorrect”) without relying on intermediate answer extraction or post-processing. This reduces engineering overhead and minimizes error propagation. Specialized verifiers show promise in supporting this end-to-end evaluation paradigm. Their structured training enables them to focus on key conclusions within noisy or verbose outputs. Enhancing robustness to formatting and expression variance is a promising direction.
@@
Specialized models can benefit from training augmentation with varied answer forms and noisy reasoning chains to boost generalization and recall. General LLMs should be guided toward more structured outputs and fine-tuned for domain-specific judgment accuracy. A hybrid pipeline is recommended: use general LLMs for high-recall coarse filtering, followed by specialized verifiers for precision filtering—balancing flexibility with accuracy.
@@ Conclusion
This benchmark presents a comprehensive evaluation of specialized verifiers and general LLMs under varying input-output constraints across multiple subjects. Results show that specialized verifiers consistently lead in accuracy, particularly in structured tasks, while larger general LLMs excel in recall and flexibility, especially when given full input and output freedom. However, both approaches have trade-offs: specialized verifiers struggle with answer diversity, and general LLMs risk misjudgment due to less precise criteria and unstructured outputs. Moreover, answer format and extraction dependency impose challenges for practical deployment. To address these gaps, future models should aim for end-to-end judgment with minimal reliance on answer extraction, while improving robustness to expression variability. We recommend enhancing generalization in specialized verifiers and guiding general LLMs toward structured outputs through targeted fine-tuning.

### v3

@@ Abstract
Large language models (LLMs) increasingly rely on reinforcement learning (RL) to enhance their reasoning capabilities through feedback. A critical challenge is verifying the consistency of model-generated responses and reference answers, since these responses are often lengthy, diverse, and nuanced. Rule-based verifiers struggle with complexity, prompting the use of model-based verifiers. However, specialized verifiers lack flexibility, while general LLM judges can be inconsistent. Existing research primarily focuses on building better verifiers, yet a systematic evaluation of different types of verifiers’ performance across domains remains lacking, severely constraining the reliable development of Reinforcement Learning with Verifiable Reward (RLVR). To address this, we propose VerifyBench–a cross-domain comprehensive benchmark for systematically evaluating verifiers. We construct about 4,000 expert-level questions covering mathematics, physics, chemistry, and biology. Each question is equipped with reference answers and diverse responses. The reliability of the evaluation is ensured through a rigorous annotation process conducted by a multidisciplinary expert team. We design a four-dimensional experimental framework to comprehensively compare the performance boundaries of specialized verifiers and general LLMs under combined conditions of extracted answers vs. complete responses, and short vs. long outputs. Our evaluation uncovers fundamental trade-offs in verifiers: while specialized verifiers achieve leading accuracy (the best model reaching 96.48% in chemistry), they exhibit deficiencies in recall; general models show stronger inclusivity but unstable precision. More importantly, we discover verifiers’ high sensitivity to input structure and inherent limitations in cross-domain generalization, providing critical insights into the bottlenecks of current verifier technology.
@@ Introduction
To fill this critical void, we introduce VerifyBench: a cross-domain comprehensive benchmark for systematic evaluation of verifiers. VerifyBench is constructed from about 4,000 expert-level questions spanning mathematics, physics, chemistry, and biology. Each question includes reliable reference answers and diverse Chain-of-Thought (CoT) (Ma et al. 2025b; Wei et al. 2022) responses. Gold-standard judgment labels are established through a rigorous, fine-grained human annotation process by a multidisciplinary expert team, ensuring unparalleled reliability. This enables systematic analysis of verifier behavior across a four-dimensional experimental framework, varying input granularity (boxed final answers vs. full reasoning traces) and output constraints.
@@
We introduce VerifyBench, a novel, multidisciplinary benchmark of about 4,000 expert-level questions across mathematics, physics, chemistry, and biology, featuring fine-grained human annotations and diverse CoT responses from state-of-the-art LLM, specifically designed for systematic verifier evaluation.
@@
We have carefully selected a total of approximately 4,000 high-quality, expert-level questions, covering four disciplines: mathematics, physics, chemistry, and biology, with around 1,000 questions per subject. These questions are designed to be challenging, incorporating rich domain-specific background, requiring advanced reasoning capabilities from LLMs, and allowing for open-ended, expressive answer spaces—thereby increasing benchmark complexity. Our definition of “expert-level” questions refers to problems typically encountered in university-level coursework or national/international academic competitions, characterized by their requirement for multi-step reasoning and synthesis of multiple concepts.
@@
Drawing from both domestic and international sources, team members manually select problems that align with our “expert-level” definition. These questions are chosen to ensure broad topical coverage, high cognitive demand, and diversity in conceptual, theoretical, and applied reasoning, always paired with correct and clear reference answers. A detailed description of our question selection criteria and source identification process is provided in the Appendix.
Each question undergoes 2 rounds of manual review. Cross-validation is conducted by graduate-level (Master’s and PhD) reviewers to identify redundant knowledge points, overly simplistic phrasing, ambiguous wording, or irrelevant content. This careful review process ensures that the questions and reference answers are professional, coherent, and free from noise. Further specifics on our review protocols and reviewer qualifications can be found in the Appendix.
@@
Differences in spelling, punctuation, and formatting are acceptable. In long-chain reasoning, correctness of logic and final answer are prioritized over surface-level language differences.
@@
This section offers a detailed summary of VerifyBench, highlighting its scale, composition, and core characteristics across disciplines. The benchmark contains about 4,000 questions, each paired with a reference answer and a model-generated response, intended to support comprehensive and rigorous verifier evaluation. To ensure balanced representation, we include about 1,000 questions from each of four major scientific domains: mathematics, physics, chemistry, and biology. Key statistics are summarized in Table 1.
@@
Total Questions | 3,989 |
@@
Total Annotated Instances | 3,989 |
@@
Verifier Mathematics Chemistry Biology Physics Overall Trained Verifier xVerify-0.5B-I 79.60%\64.92% 94.77%\96.50% 86.75%\92.86% 92.48%\94.86% 88.77%\88.66% xVerify-3B-Ia 81.28%\56.91% 94.77%\96.21% 89.56%\90.18% 91.57%\92.24% 89.24%\85.35% xVerify-8B-I 81.58%\55.80% 96.39%\96.50% 87.95%\91.07% 92.68%\92.69% 90.03%\85.47% xVerify-9B-C 83.03%\54.14% 96.28%\94.46% 89.96%\91.07% 93.78%\92.81% 90.96%\84.76% general-verify - - - - - R1-Distill-Verifier-1.5B - - - - - General LLM as Judge Qwen2.5-7B-Instruct 78.18%\91.99% 89.45%\82.22% 68.67%\83.04% 86.66%\98.17% 83.52%\93.74% Qwen2.5-14B-Instruct 78.40%\90.61% 90.05%\82.80% 54.62%\83.04% 91.28%\96.46% 84.10%\91.55% Qwen2.5-32B-Instruct 80.98%\76.24% 85.63%\61.81% 61.45%\66.96% 95.39%\97.26% 85.34%\83.58% Qwen3-8B/14B/32B - - - - - DS-R1-Distill-Qwen-7B/14B/32B - - - - -
VerifyBench emphasizes long-form, multi-step reasoning, representing a substantial challenge for existing verifiers. The high average token count for responses (Table 1) reflects the benchmark’s focus on detailed, explanatory solutions and verbose model outputs. Our two-stage human annotation pipeline ensures high-quality binary labels, with inter-annotator agreement (IAA) (Artstein 2017) scores ranging from 0.88 to 0.92, confirming annotation reliability.
@@
To simulate realistic response patterns of reasoning-capable language models, we employ QwQ-32B (Team 2025) to generate detailed CoT responses for all multi-subject questions in VerifyBench. We use the official recommended hyperparameters: temperature 0.6, TopP 0.95, MinP 0, TopK 40, and no repetition penalty. The maximum response length for QwQ-32B is set to 32,768 tokens.
@@
We evaluate two categories of LLMs as verifiers. The first category comprises general-purpose open-source LLMs, including the Qwen2.5-Instruct series (7B, 14B, 32B) (Qwen et al. 2025), the Qwen3 series (8B, 14B, 32B) (Yang et al. 2025), and the DeepSeek-R1-Distill-Qwen series (7B, 14B, 32B) (Guo et al. 2025). The second category consists of specialized verifiers trained on large-scale verification datasets, such as the xVerify series (0.5B, 3B, 8B, 9B) (Chen et al. 2025), general_verifier(Ma et al. 2025a), and R1-Distill-Verifier-1.5B (Huang et al. 2025).
Among these, models from the xVerify and Qwen2.5-Instruct series directly output binary verification decisions (e.g., “correct” or “incorrect”). In contrast, generative models—such as the Qwen3 and DeepSeek-R1-Distill-Qwen series, as well as general_verifier and R1-Distill-Verifier-1.5B—produce reasoning traces alongside final judgments.
@@
We consider the following four experimental settings, all of which include the question and reference answer as inputs:
@@
Verifier Mathematics Chemistry Biology Physics Overall Trained Verifier xVerify-0.5B-I 79.38%\54.14% 94.67%\94.46% 87.15%\91.07% 92.38%\92.81% 88.64%\84.76% xVerify-3B-Ia 81.18%\56.91% 95.03%\96.21% 88.76%\90.18% 91.88%\92.24% 89.38%\85.35% xVerify-8B-I 81.58%\55.80% 96.28%\96.50% 89.16%\91.07% 92.58%\92.69% 90.06%\85.47% xVerify-9B-C 82.78%\64.92% 96.18%\96.50% 89.96%\92.86% 93.98%\94.86% 90.86%\88.66% general-verify 68.77%\88.12% 75.13%\88.63% 73.09%\85.71% 94.32%\97.72% 79.01%\93.03% R1-Distill-Verifier-1.5B 76.18%\81.22% 80.71%\86.30% 77.91%\78.57% 88.77%\89.50% 81.91%\86.36% General LLM as Judge Qwen2.5-7B-Instruct 77.88%\47.51% 89.45%\93.59% 54.62%\29.46% 86.66%\86.64% 82.35%\75.90% Qwen2.5-14B-Instruct 78.08%\59.12% 90.15%\86.01% 62.25%\34.82% 91.67%\92.47% 84.75%\80.21% Qwen2.5-32B-Instruct 81.08%\71.27% 85.28%\96.50% 62.25%\48.21% 95.39%\97.37% 81.20%\88.36% Qwen3-8B 70.77%\74.31% 88.04%\93.88% 81.53%\83.04% 95.09%\97.37% 84.38%\90.79% Qwen3-14B 85.39%\80.11% 92.61%\95.92% 84.34%\92.86% 96.99%\98.52% 91.11%\93.68% Qwen3-32B 74.67%\70.72% 83.82%\93.00% 83.40%\91.96% 95.89%\97.15% 84.69%\90.31% DS-R1-Distill-Qwen-7B 64.66%\74.59% 75.38%\85.42% 74.80%\77.68% 91.78%\97.03% 77.07%\88.72% DS-R1-Distill-Qwen-14B 76.18%\78.73% 81.06%\85.13% 68.67%\64.29% 95.69%\97.26% 83.02%\88.66% DS-R1-Distill-Qwen-32B 72.37%\78.73% 79.10%\96.50% 72.29%\85.71% 96.59%\99.43% 81.85%\93.50%
@@
The verifier receives the full reasoning process as input but remains limited in output length. This setting challenges the model to efficiently extract conclusions from complex context. Specialized verifiers again show dominant performance, particularly in accuracy. For instance, xVerify-9B-C yields the best overall accuracy (90.96%) and leads in most subjects. While general LLMs such as Qwen2.5-7B-Instruct demonstrate high recall (e.g., 91.99% in math), they often trade precision for inclusiveness. This highlights a key contrast: general LLMs can tolerate diverse answer forms but struggle with precision. The details are shown in Table 2.
@@
From a system perspective, verifiers should ideally produce direct, structured outputs (e.g., “Correct”/“Incorrect”) without relying on the extraction of model response or verifier’s judgment result. This reduces engineering overhead and minimizes error propagation. Specialized verifiers show promise in supporting this end-to-end evaluation paradigm. Their structured training enables them to focus on key conclusions within noisy or verbose outputs. Enhancing robustness to formatting and expression variance is a promising direction.
@@
Specialized verifiers can benefit from training augmentation with varied answer forms and noisy reasoning chains to boost generalization and recall. General LLMs should be guided toward more structured outputs and fine-tuned for domain-specific judgment accuracy. A hybrid pipeline is recommended: use general LLMs for high-recall coarse filtering, followed by specialized verifiers for precision filtering.
@@ Conclusion
This benchmark presents a comprehensive evaluation of specialized verifiers and general LLMs under varying input-output constraints across multiple subjects. Results show that specialized verifiers consistently lead in accuracy, particularly in structured tasks, while general LLMs excel in recall and flexibility, especially when they have a larger model size and are given full input and output freedom. However, both approaches have trade-offs: specialized verifiers struggle with answer diversity, and general LLMs risk misjudgment due to less precise criteria and unstructured outputs. Moreover, answer format and extraction dependency impose challenges for practical deployment. To address these gaps, future models should aim for end-to-end judgment with minimal or no reliance on the extraction of model-generated answers and verifiers’ judgment results, while improving robustness to expression variability. We recommend enhancing generalization in specialized verifiers and guiding general LLMs toward structured outputs through fine-tuning.